### v1

# Back to the Basics: Rethinking Issue-Commit Linking with LLM-Assisted Retrieval

Abstract.
Issue-commit linking, which connects issues with commits that fix them, is crucial for software maintenance. Existing approaches have shown promise in automatically recovering these links. Evaluations of these techniques assess their ability to identify genuine links from plausible but false links. However, these evaluations overlook the fact that, in reality, when a repository has more commits, the presence of more plausible yet unrelated commits may interfere with the tool in differentiating the correct fix commits. To address this, we propose the Realistic Distribution Setting (RDS) and use it to construct a more realistic evaluation dataset that includes 20 open-source projects. By evaluating tools on this dataset, we observe that the performance of the state-of-the-art deep learning-based approach drops by more than half, while the traditional Information Retrieval method, VSM, outperforms it.
Inspired by these observations, we propose EasyLink, which utilizes a vector database as a modern Information Retrieval technique. To address the long-standing problem of the semantic gap between issues and commits, EasyLink leverages a large language model to rerank the commits retrieved from the database. Under our evaluation, EasyLink achieves an average Precision@1 of 75.91%, improving over the state-of-the-art by over four times. Additionally, this paper provides practical guidelines for advancing research in issue-commit link recovery.
Issue-Commit Link Recovery, Software Traceability
Jieke Shi is the corresponding author.
copyright: acmlicensedjournalyear: 2026conference: IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Software traceability involves establishing relationships between different software artifacts and is essential for safety-critical systems (Cleland-Huang et al., 2012, 2014). A critical task in this domain is issue-commit linking, which connects issues, i.e., bug reports, to the commits that resolve them (Kondo et al., 2022), playing a vital role in software provenance. It also plays a key role for developers in assessing security risks (Nguyen-Truong et al., 2022; Nguyen et al., 2022; Li et al., 2024) and gaining deeper insights about security flaws (Wu et al., 2025; Meneely et al., 2013).
Prior studies (Bachmann and Bernstein, 2009; Romo et al., 2014) have revealed that many issue-commit links can be missing during the development of large-scale projects. Manually recovering these links is not only time-consuming but error-prone, even for experienced developers (Ruan et al., 2019). To address this, several studies (Zhang et al., 2023b; Lin et al., 2021; Dong et al., 2022; Ruan et al., 2019; Sun et al., 2017a; Rath et al., 2018) have proposed learning-based approaches to automatically recover issue-commit links, achieving strong performance on datasets collected from open-source repositories. The evaluation method requires the tool to distinguish fix commits (i.e., “true links”) from plausible but non-fix commits (i.e., “false links”). However, prior studies’ evaluations often lack realism. ❶ Some studies use an unrealistic time window to select commits. For example, studies (Ruan et al., 2019; Nguyen et al., 2012; Sun et al., 2017b; Bachmann et al., 2010) rely on a narrow 7-day time window to select potential fix commits, which may miss many true links. Another example involves studies (Dong et al., 2022; Rath et al., 2018) that require the issue close time to select commits, but in practice, when commits are missing, the issue may not have a close time. ❷ Another limitation is the unrealistic false link distributions in prior evaluation datasets. Some (Lin et al., 2021; Ruan et al., 2019) use evaluation datasets where the number of false and true links is equal, an unrealistic assumption since, in reality, false links far outnumber true links (Dong et al., 2022). Recently, Zhang et al. (Zhang et al., 2023b) addressed this issue by using an imbalanced dataset with a fixed number of false links per issue. Still, this method overlooks a crucial factor: a higher commit frequency results in a larger pool of plausible commits for the tool to differentiate from the actual fix commit, making the task inherently more difficult.
Therefore, we propose to construct a more realistic evaluation dataset under the Realistic Distribution Setting (RDS). We began with an in-depth analysis of the issues and commits from open-source repositories used in prior studies (Zhang et al., 2023b; Dong et al., 2022; Lin et al., 2021; Mazrae et al., 2021). According to our observation, after fetching all commits in each repository, approximately 97% of fix commits were made within one year from issue creation. These findings suggest that, in practice, the corresponding fix commit for an issue is likely to be found among the commits made within one year after the issue’s creation. We thus include the genuinely linked commit as the true link and consider all non-fix commits made within one year of an issue’s creation as candidate commits linked to the given issue as false links when constructing evaluation datasets.
After collecting issues and commits from 20 open-source projects analyzed in prior studies (Zhang et al., 2023b; Dong et al., 2022; Lin et al., 2021; Mazrae et al., 2021), we first successfully replicate the performance of the state-of-the-art method, EALink (Zhang et al., 2023b), using the original evaluation method they reported. However, when we switched the experimental setup to the Realistic Distribution Setting, we find that the average Precision@1 of EALink dropped to 14.43%, surprisingly underperforming the Vector Space Model (VSM) (Antoniol et al., 2002), a traditional Information Retrieval (IR) technique, which achieves a Precision@1 of 46.59%, suggesting that there is great room for improving issue-commit linking using IR techniques. Meanwhile, as recent years have witnessed huge advances in the IR field, modern IR techniques that use embedding models to capture semantic-level similarities have demonstrated better performance than those traditional techniques evaluated in previous studies (Dong et al., 2022; Zhang et al., 2023b; Lin et al., 2021) that measure token-level textual similarities, such as VSM (Antoniol et al., 2002) and Latent Semantic Indexing (LSI) (Marcus et al., 2005). This also presents an opportunity to boost issue-commit linking performance and motivates our paper.
In this paper, we propose EasyLink, a novel method that integrates modern IR techniques and Large Language Models (LLMs) to enhance issue-commit linking performance in realistic settings by capturing deeper semantic relationships between issues and commits. EasyLink operates in two stages. The first stage fetches a set of commits that share similarities to a given issue, and the second stage reranks them by their relevance to the issue. Concretely, in the first stage, EasyLink leverages recent advances in Information Retrieval by using a modern vector database (optimized for high-dimensional similarity search) (Douze et al., 2024) to fetch the most similar commits for each issue. In the second stage, EasyLink prompts an LLM (GPT-4o (OpenAI, 2024)) to rerank the results. We use LLMs for their strong ability to capture the semantic relationship between issues and commits, bridging the semantic gap, as the most similar commit does not guarantee that it is the fix for the issue (Zhang et al., 2023b; Lin et al., 2021; Dong et al., 2022). EasyLink achieves an average Precision@1 of 75.91%, outperforming EALink (Zhang et al., 2023b) by 4$\times$ in our realistic datasets and by 30% in Precision@1 on EALink’s original evaluation setup. We also conduct an ablation analysis on each of EasyLink’s two stages. Our results demonstrate that advances in IR, such as context-aware dense embeddings (Wang et al., 2020; Song et al., 2020; Neelakantan et al., 2022) and efficient vector search (Douze et al., 2024; Wang et al., 2021), have been largely overlooked in the software traceability literature. Notably, even the out-of-the-box use of a vector database achieves a high average Precision@1 of 64.03%. Additionally, the strong performance of the reranking step, which improves Precision@1 by 11%, highlights the capability of LLMs to bridge the semantic gap between issues and commits (Rath et al., 2018; Guo et al., 2017). Finally, we discuss the lessons learned from our work, such as the need to consider modern IR baselines, for future research on software traceability. Our implementation, along with the evaluation dataset, has been anonymized and made available at (Package, 2025).
This paper makes the following contributions:
  1. (1)
    Constructing a more realistic evaluation dataset: To achieve a more realistic evaluation, we propose the Realistic Distribution Setting (RDS), which adjusts the number of candidate commits for generating false links according to the quantity of commits in the repository. This results in a dataset that more accurately reflects real-world practices.
  2. (2)
    A comprehensive evaluation benchmark: We include the datasets from recent studies (Zhang et al., 2023b; Dong et al., 2022; Lin et al., 2021; Mazrae et al., 2021). Our benchmark includes 9,319 issues from 20 projects, with an average of 1,530 false links constructed per issue. To the best of our knowledge, this dataset is the largest in the literature.
  3. (3)
    Reevaluation of the state-of-the-art approach: After successfully replicating the strong performance of EALink (Zhang et al., 2023b), the state-of-the-art approach, we switched the evaluation procedure under RDS, which offers a greater challenge due to a higher number of false links. On the same set of projects used in the evaluation of EALink, its average Precision@1 of 53.67% decreases to 28.05%, underperforming traditional IR baselines.
  4. (4)
    A new state-of-the-art for issue-commit linking: We propose EasyLink, which leverages modern IR techniques, including an off-the-shelf database FAISS (Douze et al., 2024), and addresses the problem of semantic gap (Rath et al., 2018; Guo et al., 2017) by prompting an LLM to rerank the retrieved results. In our evaluation, EasyLink improves over EALink in average Precision@1 from 14% to 76%.
The rest of the paper is organized as follows. Section 2 outlines the background of this work, including the limitations of existing work. Section 3 details the evaluation dataset construction under the Realistic Distribution Setting (RDS). Section 4 introduces EasyLinkSection 5 describes the experimental setup. Section 6 presents the experimental results. Section 7 discusses the lessons learned and threats to validity. Section 8 reviews related work.  Section 9 concludes the paper. Finally,  Section 10 presents the details of data availability.
2. Background
2.1. Issue-Commit Linking Recovery
An example of an issue, the incorrectly top-ranked commit, and the correct commit. The input issue, with summary and description, is shown below. Issue ID: NETBEANS-803 Summary: nb-javac 11 upgrade in NetBeans Description: Should cover below tasks in NetBeans - nb-javac 11 testing : Run tests for modules java.completion, java.editor, java.editor.base, java.hints, java.source, java.source.base, lib.nbjavac Update libs.javacimpl and libs.javacapi jars, upload updated nb-javac jars Upload nb-javac module jars in update center   The top-ranked commit message in the initial results is shown below but is incorrect. Commit ID: 4fd115aeae3b8423de9ced22d52914e60a1c5800. Updation for external nb-javac jar in libs.javacapi and libs.javaimpl modules with nb-javac jar for jdk-12   The correct commit message is shown below. Commit ID: 3055661e4dd1c7d587012c917cbec31b27ae9e34 Uptake nb-javac 11 jars for java tests runtime
Figure 1. Example of an issue with the incorrect top-ranked commit and the correct commit. An issue-commit linking approach has to bridge the semantic gap and distinguish the correct commit from similar ones in a potentially large set.
Issue-commit links play an essential role in maintaining software traceability, supporting critical tasks such as impact analysis (Aung et al., 2020; Hindle et al., 2008; Purushothaman and Perry, 2005), regression testing (Naslavsky and Richardson, 2007), and project management (Panis, 2010). Due to the high cost of manually maintaining these links, they are often incomplete (Izadi et al., 2022), highlighting the need for automated methods. Automatic issue-commit linking considers a large set of commits that are made in the period after each issue’s creation and identifies the right commits that address the issue. Figure 1 shows an example that demonstrates the challenge of automatically linking issues to commits. It presents an issue along with two commits. The first commit shares more matched words and has the highest similarity to the issue, but it is not the fix, while the second commit is the correct fix commit. An issue-commit linker that considers keyword counts would incorrectly prioritize the first commit as there are a greater number of matches of the keyword nb-javac. While both commits mention nb-javac, the top-ranked commit updates the JAR for JDK-12, whereas the issue specifies that nb-javac 11 should be upgraded, tested, and deployed. This highlights two challenges. First, it shows the need for approaches that go beyond surface-level similarities. Second, it demonstrates the sensitivity of the evaluations of their effectiveness to the number of commits that share keywords or resemble the ground-truth commit.
Many automatic issue-commit link recovery methods have been proposed. Some studies employ traditional feature- and rule-based methods  (Schermann et al., 2015; Wu et al., 2011; Nguyen et al., 2012; Bachmann et al., 2010), which rely on predefined heuristics such as keyword matching or the recency of the commits. However, these heuristics tend to be inadequate  (Bird et al., 2009). Some approaches (Mazrae et al., 2021; Rath et al., 2018; Sun et al., 2017a; Le et al., 2015; Sun et al., 2017b) adopt traditional machine learning techniques, such as support vector machines, to reduce the reliance on manual rules. Recently, deep learning methods  (Ruan et al., 2019; Xie et al., 2019; Guo et al., 2017) have demonstrated improved performance. A key challenge of the task is the semantic gap that exists between issues and commits (Zhang et al., 2023b; Lin et al., 2021; Dong et al., 2022). Prior work (Lin et al., 2021) attempted to address this issue using BERT-based methods, such as CodeBERT (Feng et al., 2020), which leverage contextual understanding. The state-of-the-art method, EALink (Zhang et al., 2023b), employs knowledge distillation to transfer knowledge to a smaller model and utilizes multi-task learning to improve both accuracy and efficiency.
2.2. Limitations in Evaluation
The experiments in prior studies evaluate approaches based on their ability to distinguish the true link from false links constructed using other commits. However, these studies use unrealistic methods to construct false links and the evaluation dataset.
Unrealistic Time Window Selection: Some works (Ruan et al., 2019; Nguyen et al., 2012; Sun et al., 2017b; Bachmann et al., 2010) assume that fix commits fall within a 7-day time window before or after the issue’s create/update/close time or the comment create time, treating all other commits in this period as non-fix commits. However, this approach is inadequate when fixes take a longer time, requiring tools to distinguish the correct fix commit from more plausible commits. From our preliminary analysis of selected projects, we found that only 59% of issues had a corresponding fix commit within seven days of their creation. Additionally, some works (Dong et al., 2022; Rath et al., 2018) require the issue’s close time to select commits. Under a practical scenario, the issue fix/close date is unknown—precisely also why issue-commit recovery is necessary—and approaches need to distinguish the fix commit while the issue remains open. Therefore, selecting commits based on the issue’s fix/close date may also lack realism.
Unrealistic False Link Distribution: When constructing the evaluation dataset, some works (Ruan et al., 2019; Lin et al., 2021; Mazrae et al., 2021) use a balanced dataset (i.e., an equal number of false links and true links) to evaluate their tools. This evaluation setting is unrealistic because, in reality, false links outnumber true links (Dong et al., 2022). Zhang et al. (Zhang et al., 2023b) address the balanced dataset limitation by constructing a fixed number of 99 false links per issue. While this results in more false links than true links, it still overlooks that the number of potentially linked commits depends on the number of commits made in the same time period as the ground-truth link. Moreover, as shown in Figure 2, when sampling unrelated commits to construct the false links, prior work (Zhang et al., 2023b) only selects commits from other ground-truth links. In other words, they construct false links for an issue by connecting it only to other commits that are already linked to other issues. In practice, an approach has to consider all commits from the same time period, regardless of whether they are linked to specific issues. This is a superset of commits compared to the ground-truth links. In Section 3, we show that the number of commits made within a one-year time frame is at least 80% larger than the constant number of 99 commits considered by Zhang et al. (Zhang et al., 2023a). As a result, the evaluation setups of prior work may not adequately reflect how these approaches would be used in practice.
Figure 2. Illustration of the evaluation limitation in prior work. $C_{1}$ and $C_{2}$ are the fix commits of issue $I_{1}$. False links will be constructed for $I_{1}$. $C_{3}$ and $C_{4}$ are commits already linked to the issue in the true links dataset. $C_{a}$, $C_{b}$, $C_{c}$, and $C_{d}$ are commits present in the repository but not in the true links dataset, and they are ensured not to be the fix commit of $I_{1}$.
3. Evaluation Dataset Construction Under the Realistic Distribution Setting
This section describes our new evaluation setting, Realistic Distribution Setting (RDS), for constructing a more realistic evaluation dataset. By adaptively constructing false links for an issue based on the commits in the period after each issue creation, aligning with the repository’s development activity, our method addresses the limitations of prior evaluations.
We combine datasets from four recent studies (Zhang et al., 2023b; Lin et al., 2021; Dong et al., 2022; Mazrae et al., 2021) into a unified benchmark covering 20 projects, ensuring broader coverage and a consistent benchmark for comparison. Evaluations in the previous studies (Zhang et al., 2023b; Lin et al., 2021; Mazrae et al., 2021) did not use a shared benchmark, and each of their datasets had only up to 12 projects (Mazrae et al., 2021). Our dataset fills the need for a large, shared benchmark.
3.1. Ground Truth Dataset Selection
We selected datasets from four recent studies, covering 20 open-source software projects over a span of 20 years. We include the same issues and true links from these datasets, resulting in a total of 9,319 issues in the benchmark. For datasets with incomplete information detected through a manual check, we re-fetched the necessary data, such as issue comments, code diffs, and committed files (i.e., the source code files after applying the commits).
  • From Zhang et al.’s dataset (Zhang et al., 2023b), we include the issues from all six projects, Ambari, Calcite, Groovy, Ignite, Isis, and Netbeans, in our benchmark.
  • From Dong et al.’s dataset (Dong et al., 2022), we include the issues from five projects, Pig, Maven, Infinispan, Drools, and Derby. This dataset also includes Groovy, which is already included in our benchmark from Zhang et al.’s dataset (Zhang et al., 2023b). As they were missing issue comments, code diffs, and committed files, we refetched these data.
  • From Lin et al.’s dataset (Lin et al., 2021), we include the issues from all three projects: Pgcli, Flask, and Keras. As the issue comments and committed files were missing, we refetched them.
  • From Mazrae et al.’s dataset (Mazrae et al., 2021), we include the issues from six projects, Beam, Flink, Freemarker, Airflow, Arrow, and Cassandra, out of 12 projects. The six other projects from this dataset are already included in Zhang et al.’s dataset (Zhang et al., 2023b). As this dataset provides only postprocessed data, we re-fetched the original issue summaries, issue descriptions, issue comments, commit messages, code diffs, and committed files.
3.2. Preparing Commits
For constructing false links, the set of all commits in each repository is required. First, we clone the GitHub repository locally, which enables faster access to commit data without API requests. From the cloned repository, all commit IDs are obtained using the git log command. Each commit is then processed to retrieve its metadata, including the parent commit IDs, author, committer, commit time, and commit message, using git show. Additionally, the list of modified file paths for each commit is identified, and their content at the specific commit state is retrieved. The corresponding code diffs are also extracted and stored. Table 1 presents the number of commits fetched for each project in the column “#Commits”.
The following information is fetched during the process:
  • Commit ID: A unique hash value assigned to each commit, serving as its identifier.
  • Parent Commit IDs: The hash(es) of the immediate predecessor commit(s) of the current commit.
  • Author: The person who originally wrote the changes.
  • Committer: The person who applied the changes to the repository.
  • Commit Time: The commit’s creation timestamp.
  • Commit Message: The description of the commit.
  • Changed Files: List of paths of the files modified in the commit.
  • Code Diffs: The changes introduced by the commit.
  • Committed Files: Source code files after applying the commit.
Note that while our work only requires the commit ID, commit time, and commit message, we collect all commit information since other approaches may utilize this additional data.
Figure 3. Overview of EasyLink. EasyLink consists of two key steps—the first step utilizes a vector database to retrieve initial ranked results, and the second step prompts an LLM to rerank the results.
3.3. Constructing False Links
1
Input: test_set_true_links
  $\triangleright$ Contains true issue-commit pairs labeled as 1
2
Input: commits_pool
  $\triangleright$ Contains all fetched commits (Subsection 3.2)
3
Result: evaluation_dataset
4
5 Function EvalSetGen(test_set_true_links, commits_pool):
6       evaluation_dataset $\leftarrow$ $\emptyset$
7       sampled_issue_id_list $\leftarrow$ test_set_true_links[issue_id].unique()[:1000]
8
9      for each issue_id in sampled_issue_id_list do
10             true_links $\leftarrow$ test_set_true_links[issue_id]
11             issue $\leftarrow$ true_links[0].issue_info
12
[…]
14                 commit.time $\geq$ issue.create_time and
15                 commit.time $\leq$ issue.create_time + $\epsilon$   and
16                 commit $\neq$ issue.fix_commit }
17
[…]
19                 commit $\in$ candidate_commits }
20
21            evaluation_dataset.append(false_links)
22             evaluation_dataset.append(true_links)
23
24
25      return evaluation_dataset
Algorithm 1 False Links Construction under RDS
For a more realistic evaluation dataset, the number of false links should match the actual number of commits that may be viable for linking to each issue in practice. For each issue in the evaluation dataset, under the Realistic Distribution Setting (RDS), false links will be adaptively constructed to better align with the repository’s level of activity. We include all commits submitted in the time window within which the true commit can appear. To determine the size of the time window, we perform an analysis and find that approximately 97% of all issues have their ground truth commit submitted within one year of their creation (7-day time window: 59%, 30-day time window: 77%, 6-month time window: 92%). This observation aligns with previous studies analyzing bug reports (e.g., Rodrigues et al. (Rodrigues et al., 2020), Zhang et al. (Zhang et al., 2023a)). Consequently, we apply a one-year time window to select candidate commits from the commits pool (as described in Subsection 3.2) for each issue. Specifically, for a true link denoted as $t_{i}=\{I_{i},C_{i}\}$, where $I_{i}$ represents the issue and $C_{i}$ the commit in the true link, the candidate commit $C_{j}$ is determined using Equation 1. If the commit time of $C_{j}$ is within one year after the creation time of $I_{i}$, we treat $C_{j}$ as a candidate commit. By linking $C_{j}$ to $I_{i}$, we generate the false link $f_{i}=\{I_{i},C_{j}\}$.
[…]
$\land\;committed(C_{j})\leq created(I_{i})+\epsilon,$
[…]
Table 1. Statistics of the benchmark
Project name #Commits #Unique issue_id Average # of false links per issue
Ambari 24809 1000 3978.26
Calcite 5889 551 516.31
Groovy 20862 1000 1064.79
Ignite 28869 1000 2611.81
Isis 24945 652 1265.91
Netbeans 10502 159 1279.68
Derby 8040 43 363.93
Drools 16585 182 630.74
Infinispan 17078 399 1110.68
Maven 14685 61 377.15
Pig 3675 45 197.44
Flask 5353 151 357.88
Keras 11248 111 454.23
Pgcli 2364 105 356.92
Airflow§ 27123 961 1882.61
Arrow§ 16947 1000 2058.99
Beam§ 43595 865 6096.72
Cassandra§ 29867 25 2560.84
Flink§ 35767 1000 3251.68
Freemarker§ 2492 9 179.00
Note: Each project is annotated with a superscript representing its original dataset source: Zhang et al. (Zhang et al., 2023b), Dong et al. (Dong et al., 2022), Lin et al. (Lin et al., 2021), § Mazrae et al. (Mazrae et al., 2021)
Table 1 shows the statistics of the evaluation dataset. The dataset includes the same issues as used in the evaluation of prior studies (Zhang et al., 2023b; Dong et al., 2022; Lin et al., 2021; Mazrae et al., 2021). Each issue has an average of 1,530 false links, with a minimum of 179 (Freemarker). In contrast, for each issue, EALink (Zhang et al., 2023b) constructs a constant number of 99 false links per issue, which is substantially smaller than the number of commits within a one-year time frame from issue creation. Note that an issue may be fixed by one or more commits (Zhang et al., 2023b). In our evaluation dataset, 17% of issues were fixed by multiple commits.
4. The EasyLink Approach
This section details EasyLinkFigure 3 shows an overview of EasyLink, which consists of two stages: the first stage uses a vector database for scalable retrieval to fetch commits ranked by their similarity to the issue and the second stage prompts an LLM to rerank them by their relevance to the issue. We elaborate the key steps, vector database indexing and retrieval, and LLM-assisted reranking below.
4.1. Vector Database Indexing and Retrieval
The process begins by embedding the commit messages using an embedding tool, retaining the commit ID as metadata. Similarly, the issue summary and description are concatenated and embedded together, with the issue ID retained as metadata. To generate the embeddings, we utilize the Sentence-Transformers library (Reimers and Gurevych, 2019) with the all-MiniLM-L6-v2 model, a lightweight and efficient transformer based on Microsoft’s MiniLM architecture (Wang et al., 2020). We chose this model because it is the most downloaded and most liked sentence-similarity model on Hugging Face, indicating strong community trust and widespread adoption. This model employs self-attention distillation to capture contextual information effectively while maintaining computational efficiency. Given an input text $T$, the embedding process can be defined as follows:
(2) $\mathbf{E}=\text{MiniLM}(T)$
where $\mathbf{E}\in\mathbb{R}^{384}$ represents the output embedding vector in a 384-dimensional space.
After generating embeddings, we use a vector database for indexing and retrieval. We use FAISS (Douze et al., 2024)—an off-the-shelf library designed for efficient nearest neighbor (NN) retrieval in high-dimensional spaces. We believe that switching FAISS with another vector database with comparable capability would produce similar results, as shown in  Subsection 6.3,
To retrieve commits from the vector database, we compute the squared L2 (Euclidean) distance between the query embedding and the stored embeddings, as we use the vector database through LangChain (Chase, 2022), where L2 distance is used by default. The computation is given by the following equation:
(3) $distance(\mathbf{q},\mathbf{e}_{i})=|\mathbf{q}-\mathbf{e}_{i}|_{2}^{2}$
where $e_{i}$ is the element of the commit message embeddings $\mathbf{E}=\langle\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{n}\rangle$ that are indexed by the vector database for efficient access, and $\mathbf{q}$ is an issue embedding. A shorter distance between the query issue and the commit message indicates greater similarity, enabling the ranking of commit messages and the generation of a list of candidate commits.
4.2. LLM-Assisted Reranking
Due to the semantic gap between issues and commits (Zhang et al., 2023b; Lin et al., 2021; Dong et al., 2022), retrieving the most similar commit does not guarantee finding the one that fixed the issue. While the correct commit may have been retrieved, it could be obscured by incorrect commits that exhibit greater similarity to the issue. To address this, we rerank the retrieved commits using a large language model (LLM).
In this phase, the top-$k$ commits from the initial retrieval are provided to the LLM with a structured prompt, as shown in Figure 3. The prompt includes the issue text (summary and description), along with each commit’s ID and message. We also specify an output format instructing the LLM to return a reranked list of commit IDs. The LLM will analyze these commits and produce a reranked list based on contextual understanding and issue-commit relevance. Outputs that do not match the expected format (0.1% of the time) default to the initial retrieval results.
In detail, we adopt a zero-shot approach guided by prompts with ChatGPT (gpt-4o), following recent studies (Wang et al., 2024; Islam and Moushi, 2024; Fuchß et al., 2025), which demonstrate gpt-4o advanced capabilities in understanding complex textual relationships. The parameter $k$ controls the additional cost incurred for increasing precision. While it is possible to rerank all fetched commits for higher precision, this requires a longer prompt for the LLM, requiring more computation resources. For the value of $k$, we select $k$=10 for its balance of precision and efficiency. We later show that this allows a high Precision@1 without incurring a high cost. This will be discussed in Subsection 6.3.
5. Experimental Setup
5.1. Research Questions
This work aims to answer the following research questions (RQs):
RQ1: How does the state-of-the-art tool perform on a realistic evaluation dataset? This question investigates the performance of the state-of-the-art EALink (Zhang et al., 2023b) on the evaluation dataset constructed under the Realistic Distribution Setting (RDS). First, we replicate the successful performance of EALink on their original evaluation dataset. Next, we investigate the performance of EALink after expanding the evaluation onto a larger benchmark. Afterwards, to understand the sensitivity of its performance to the evaluation dataset, we investigate how much the performance of EALink changes when evaluated under Realistic Distribution Setting and compare it with the traditional IR method, VSM.
RQ2: How does EasyLink perform on the same realistic evaluation dataset? This research question is concerned with the effectiveness of EasyLink. We assess EasyLink in terms of both its ability to distinguish true links from false links and its efficiency. We analyze the sensitivity of the performance of EasyLink to the evaluation setup by comparing its performance on the different evaluation methods.
RQ3: Does EasyLink’s performance change under different configurations? This question aims to investigate the effect of different settings of EasyLink. We explore different embedding models for the vector database and various $k$ value settings for the LLM-assisted stage to assess whether these changes will significantly affect EasyLink’s performance.
5.2. Experiment Setting
5.2.1. Hardware Configuration
The experiments were conducted on a machine equipped with two Intel(R) Xeon(R) Platinum 8480C CPUs @ 3.80GHz, 2.0 TiB of main memory, and one NVIDIA H100 80GB HBM3 GPU.
5.2.2. LLM Setup
We utilized the ChatGPT model GPT-4o provided by OpenAI, specifically the gpt-4o-2024-11-20 version, with its default configuration. The temperature was set to the default value of 1.0, and the model was sampled once per query. The input token size was limited to a 128k token context window per API constraints, with no additional adjustments or fine-tuning.
5.2.3. Baseline Description
We use EALink [^1] We use EALink provided code and data from https://github.com/KDEGroup/EALink. (Zhang et al., 2023b) as the baseline, which is a state-of-the-art tool that outperforms T-BERT (Lin et al., 2021) and DeepLink (Ruan et al., 2019). It distills knowledge from CodeBERT (Feng et al., 2020) into a smaller model, fine-tuned with multi-task contrastive learning. We follow its original methodology, using the provided code and the same hyperparameters. As a fundamental baseline, we also run VSM [^2] We implemented VSM using the Gensim library (https://pypi.org/project/gensim) (Salton et al., 1975), which is widely used in other studies (Zhang et al., 2023b; Lin et al., 2021; Dong et al., 2022).
5.3. Evaluation Metrics
We adopt the same metrics used by Zhang et al. (Zhang et al., 2023b), which use standard metrics for information retrieval tasks (Radlinski and Craswell, 2010; Sakai and Kando, 2008): Precision@$k$ (P@$k$), Normalized Discounted Cumulative Gain (NDCG@$k$), Mean Reciprocal Rank (MRR), and Hit Ratio (Hit@$k$) for evaluation. We also include Recall@$k$, which was overlooked in prior studies.
  • Precision@k evaluates the proportion of relevant commits (i.e., commits that belong to the correct issue-commit links) within the top $k$ retrieved results:
[…]
    where $Q$ is the query set, $|Q|$ its size, and $\text{Rel}_{i}$ the number of correctly linked commits in the top $k$ results for query $i$.
  • Hit@k measures the likelihood that at least one correct commit appears within the top $k$ retrieved results:
[…]
    where $\mathbb{I}(\cdot)$ returns 1 if the highest-ranked relevant commit for query $i$ is within the top $k$, and 0 otherwise. Hit@1 is equivalent to Precision@1.
  • Recall@k evaluates the proportion of relevant commits retrieved within the top $k$ results:
[…]
    where $Q$ is the query set, $|Q|$ its size, $\text{Rel}_{i}$ the number of retrieved relevant commits, and $\text{TotalRel}_{i}$ the total relevant commits for query $i$.
  • MRR (Mean Reciprocal Rank) evaluates how early the first relevant commit appears in the ranked list for each query:
    (7) $\text{MRR}=\frac{1}{|Q|}\sum_{i=1}^{|Q|}\frac{1}{\text{Rank}_{i}},$
    where $|Q|$ is the number of queries, and $\text{Rank}_{i}$ is the position of the first correctly linked commit for query $i$. A higher MRR indicates earlier retrieval of relevant commits.
  • NDCG@k (Normalized Discounted Cumulative Gain) assesses how well relevant commits are ranked within the top $k$ retrieved results:
[…]
    where $Z_{k}$ is a normalization factor ensuring the ideal ranking achieves a value of 1. The term $r_{i}$ represents the relevance score of the commit at position $i$ ($r_{i}=1$ for a correct match, 0 otherwise).
6. Results
6.1. RQ1: How does the state-of-the-art tool perform on a realistic evaluation dataset?
To ensure we replicate EALink (Zhang et al., 2023b) correctly and perform a fair comparison, we ran EALink on both the evaluation dataset constructed using its original method and the dataset created under the Realistic Distribution Setting (RDS), then compared the results. In the original experiments of Zhang et al., EALink was trained using a balanced dataset and then tested on an imbalanced dataset. We reused the same code for training.
Table 2. Performance of EALink on evaluation datasets constructed using different methods
False Links=99 RDS
P@1 (Hit@1) 32.52 14.43 ($\downarrow 55.63\%$)
P@10 7.08 3.64 ($\downarrow 48.59\%$)
Hit@10 59.46 30.76 ($\downarrow 48.27\%$)
Recall@10 56.68 28.06 ($\downarrow 50.49\%$)
MRR 41.16 20.21 ($\downarrow 50.90\%$)
NDCG@1 20.52 9.10 ($\downarrow 55.65\%$)
NDCG@10 30.84 15.52 ($\downarrow 49.67\%$)
Note: Results are averaged over 20 projects. The column False Links=99 shows EALink’s performance on a dataset constructed with a fixed 99 false links per issue, while the column RDS shows its performance on a dataset constructed under Realistic Distribution Setting (RDS). The parentheses in the RDS column show the percentage change in performance.
Following EALink’s original evaluation (Zhang et al., 2023b), we first used their false link generation method to construct an evaluation dataset using projects provided by them: Ambari, Calcite, Groovy, Ignite, Isis, and NetBeans. For any given issue, it randomly samples 99 of the commits in the ground truth test dataset to construct false links. Next, with this evaluation dataset, which we refer to as the original evaluation dataset, we evaluated EALink. EALink obtains an average Precision@1 of 53.67%, which matches the 53.90% Precision@1 reported in the paper, indicating that our replication was successful. Then, we expanded the evaluation to 20 projects (introduced in  Subsection 3.1). The results, shown in the first column of Table 2, indicate an average Precision@1 of 32.52%.
Then, we ran EALink on the evaluation dataset constructed under the Realistic Distribution Setting. Unlike the original evaluation dataset of Zhang et al. (Zhang et al., 2023b), which had a fixed number of 99 false links per issue, our evaluation generates more false links for repositories with a larger number of commits. This results in an average number of false links per issue of 1529.78. In the second column of Table 2, we present the results of running the tool on our evaluation dataset, which shows a significant performance drop compared to the original evaluation dataset. Across the 20 projects, EALink’s Precision@1 decreases to 14.43%, a decline of 55.63%. Similarly, Hit@10 drops from nearly 60% to about 30%, reducing by half. On the six projects provided by EALink (including Ambari, Calcite, and four others), the average Precision@1 decreases from 53.67% to 28.05%. Additionally, we evaluated the traditional IR method, VSM, on the same evaluation dataset. Its results are presented in Table 4. Surprisingly, EALink underperforms the traditional IR method, VSM. VSM achieves a Precision@1 of 46.59% and demonstrates better performance across all other metrics. These results suggest that the use of VSM would be preferred over EALink in a realistic setting.
Answer to RQ1: We constructed an evaluation dataset under the Realistic Distribution Setting (RDS), which considers more false links per issue when there is higher commit activity during a given period, better reflecting the practical use of an issue-commit link recovery technique on a repository. Under this evaluation, the average Precision@1 of the state-of-the-art tool declines to 14.43% compared to a VSM baseline with a Precision@1 of 46.59%.
6.2. RQ2: How does EasyLink perform on the same realistic evaluation dataset?
6.2.1. Comparison of Effectiveness
Table 3. Comparison of linking effectiveness on the original dataset with a constant 99 false links constructed per issue. This comprises the six projects provided by Zhang et al. (Zhang et al., 2023b) (Ambari, Calcite, Groovy, Ignite, Isis, and NetBeans)
EALink Vector DB EasyLink
P@1 (Hit@1) 53.67 81.59 89.58($\uparrow 66.90\%$)
P@10 8.94 11.08 11.08($\uparrow 23.93\%$)
Hit@10 72.56 92.44 92.44($\uparrow 27.39\%$)
Recall@10 69.03 89.43 89.43($\uparrow 29.55\%$)
MRR 60.13 85.57 90.80($\uparrow 51.00\%$)
NDCG@1 33.86 51.48 56.52($\uparrow 66.92\%$)
NDCG@10 41.33 56.07 58.03($\uparrow 40.41\%$)
Note: The parentheses in the EasyLink column show the improvements over EALink. Bold numbers indicate the highest performance. All values are in %.
Table 4. Comparison of linking effectiveness on the dataset constructed under RDS
Metric EALink VSM Vector DB EasyLink
P@1 (Hit@1) 14.43 46.59 64.03 75.91 ($\uparrow 426.1\%$)
P@10 3.64 8.70 10.82 10.82 ($\uparrow 197.3\%$)
Hit@10 30.76 70.94 85.35 85.35 ($\uparrow 177.5\%$)
Recall@10 28.06 67.22 81.48 81.48 ($\uparrow 190.4\%$)
MRR 20.21 55.10 71.59 79.92 ($\uparrow 295.4\%$)
NDCG@1 9.10 29.39 40.40 47.89 ($\uparrow 426.3\%$)
NDCG@10 15.52 39.45 49.37 52.34 ($\uparrow 237.2\%$)
Note: Results are averaged over 20 projects (%). The parentheses show the improvements over EALink. Since P@10, Hit@10, and Recall@10 only assess whether the true commits are included within the top 10 results, without considering its exact rank, they are the same for both the vector database and EasyLink. Bold numbers indicate the highest performance.
To ensure a fair comparison, we first evaluated the vector database and EasyLink on the dataset constructed using the EALink method (Zhang et al., 2023b), which generates 99 false links per issue and includes the six projects provided by EALink. The results, shown in Table 3, indicate that EasyLink achieves a Precision@1 of 89.58%, significantly outperforming EALink’s 53.67%. The vector database method also outperforms EALink, achieving a Precision@1 of 81.59%.
Furthermore, we compared them using our more realistic evaluation dataset. As presented in Table 4, the vector database approach achieves an average Precision@1 of 64.03%, a significant improvement over EALink’s 14.43%. With LLM-assisted reranking, EasyLink further enhances Precision@1 by an average of 11.88%, reaching 75.91%, representing a 426.1% increase compared to EALink. Following the Mann-Whitney U test (Mann and Whitney, 1947), the improvement of EasyLink over EALink in every metric is statistically significant (p-value $<$ 0.01) and exhibits a large effect size (Cohen, 2013) (Cohen’s D $>$ 0.8). These results demonstrate that EasyLink is effective in issue-commit link recovery and that LLMs can be effectively leveraged to enhance performance.
6.2.2. Comparison of Efficiency
Table 5. Comparison of training and testing time cost
EALink Vector DB EasyLink
Train (Total / Per Link) 100.68h / 4.39s N/A N/A
Test (Total / Per Issue) 17.78h / 6.87s 2.60h / 1.01s 17.76h / 6.86s
Note: “Total” indicates the overall time for all 20 projects. “Per Link” (training) is the average time per link, computed as total training time divided by the number of training links. “Per Issue” (testing) is the average time per issue, based on total testing time divided by the number of evaluation issues. Testing times for Vector DB and EasyLink include embedding, indexing, and retrieval.
Efficiency is an important aspect of software traceability, especially when used in large-scale industrial settings (Cleland-Huang et al., 2012; Spanoudakis and Zisman, 2005). Therefore, we calculated the training and testing times required for three approaches—EALink, the vector database, and the EasyLink—to compare their efficiency. The results, shown in Table 5, indicate that for the 20 projects, EALink requires 100.68 hours for training, with an average of 4.39 seconds per link in the training set. However, approaches using a vector database do not require training on a specific issue-commit link dataset, making them easier to use and more practical with less human effort.
Regarding testing time, EALink requires a total of 17.78 hours for the 20 projects, averaging 6.87 seconds per issue, while the vector database approach takes only 2.60 hours in total, averaging 1.01 seconds per issue. These differences are statistically significant according to the Mann-Whitney U test (Mann and Whitney, 1947) (p-value $<$ 0.01) and demonstrate a large effect size (Cohen’s D $>$ 0.8). Additionally, the LLM-assisted method (i.e., EasyLink) takes 17.76 hours for testing, which is slightly faster than EALink but achieves a significantly higher performance. These results demonstrate that the use of vector database is an efficient solution. It eliminates the need for pretraining models and reduces testing time. By incorporating an LLM reranking step, EasyLink achieves a high precision while eliminating the need for training models.
Answer to RQ2: Our results show that on the realistic evaluation dataset, EasyLink achieves a Precision@1 of 75.91%, outperforming EALink (which achieves 14.43%) by 426.1%. Moreover, using the vector database-based method eliminates the need for pretraining on task-specific data, making it a more efficient solution.
6.3. RQ3: Does EasyLink’s performance change under different configurations?
6.3.1. Vector Database: Evaluating Different Embedding Models
We evaluated two additional embedding models for use in the vector database. The first model, all-mpnet-base-v2, built upon Microsoft’s MPNet architecture (Song et al., 2020), is a sentence transformer and the second-most-downloaded sentence-similarity model on Hugging Face (the most popular is used in EasyLink). The second model, text-embedding-ada-002, is the default model for OpenAI embeddings (Neelakantan et al., 2022), which is a larger model designed for higher precision.
Table 6. Comparison of linking effectiveness while varying the choice of embedding method
MiniLM MPNet OpenAI
P@1 (Hit@1) 64.03 61.22 72.16
P@10 10.82 10.37 11.09
Hit@10 85.35 82.39 88.64
Recall@10 81.48 78.58 84.73
MRR 71.59 69.02 78.40
NDCG@1 40.40 38.65 45.53
NDCG@10 49.37 47.62 52.67
Test Time (hour) 2.60h 3.10h 29.61h
Model Size 22.7M params 109M params Cloud-based
Note: Results are averaged over 20 projects. All values are in %, except for test time and model size.
Table 6 presents the results comparing three embedding models. The comparison between MiniLM (used in EasyLink) and MPNet shows that the model we selected in our tool requires less time while achieving relatively better performance. Although OpenAI embeddings provide better performance—with Precision@1 increasing from 64.03% to 72.16% (an improvement of 8.13%)—they require 29.61 hours to complete the experiment, approximately 11 times longer than the MiniLM model. Given the high computational cost, MiniLM is a more practical option.
6.3.2. LLM-Assisted Reranking: Exploring Different Top-$k$ Values
To further analyze the effect of the $k$ setting, we conduct experiments with different top-$k$ values, specifically $k=5,10,15,20$. We conducted this experiment across 20 projects using the realistic evaluation dataset. The results in Figure 4 show that increasing $k$—meaning reranking a larger set of top results—generally leads to better performance, specifically higher Precision@1 and NDCG@1. However, this improvement comes with a significant increase in time cost and requires processing more input tokens, leading to a higher cost in invoking OpenAI API. While setting $k=5$ results in relatively lower performance, increasing $k$ to 10, 15, or 20 does not lead to significant performance gains. Considering performance, time, and resource costs, we find that $k=10$ provides the best trade-off. Additionally, compared to the straight line in the figure representing the Precision@1 of EALink’s when run on the same evaluation dataset, our approach outperforms EALink across different $k$ values. This shows that its performance does not rely on tuning $k$.
Answer to RQ3: For vector database retrieval, using different embedding models with similar runtime costs yields comparable results. In the LLM-assisted reranking stage, increasing the $k$ value consistently improves performance over the baseline but increases test time cost.
7. Discussion
7.1. Lessons learned and Implications
Keep your feet on the ground – evaluation should match practice. Our results suggest that the performance of issue-commit linking approaches is sensitive to their evaluation setups, which emphasizes the importance of the evaluation of proposed techniques to reflect the conditions under real-world practices. Future research should use evaluations that better reflect a practical usage scenario.
Figure 4. Effect of varying $k$ in EasyLink during reranking: A higher $k$ slightly raises performance (left axis: P@1, MRR, NDCG@1) while significantly increasing test time cost (right axis).
Back to the basics – IR techniques are strong baselines. Our results showed that the VSM baseline, a dated IR model (Antoniol et al., 2002) proposed by Wu et al. (Wu et al., 2011) for issue-commit linking, was effective. This is consistent with the “Easy over hard” principle advocated by Fu and Menzies (Fu and Menzies, 2017). This finding has practical implications – practitioners may find that the use of simple and fast approaches match the performance of more complex approaches.
Don’t forget your roots – updating baselines. Studies on issue-commit linking have included traditional IR baselines, such as the Vector Space Model (VSM) (Salton et al., 1975), Latent Dirichlet Allocation (LDA) (Blei et al., 2003), and Latent Semantic Indexing (LSI) (Deerwester et al., 1990). However, recent studies continue to rely on traditional methods as baselines without considering recent advances in the IR literature. In our experiments, the out-of-the-box use of a modern baseline, the vector database FAISS (Douze et al., 2024), already achieves a strong performance. This shows that vector databases provide an effective retrieval approach for issue-commit linking, highlighting the need to update baselines to reflect the latest advancements in IR methods. We call for the need for newly proposed methods to be evaluated with baselines that are continuously updated to accurately measure real progress.
From good to great – effectiveness of LLMs in retrieval refinement. Our experiments demonstrate that a reranking step using LLMs is highly effective in overcoming the semantic gap, consistent with findings in information retrieval systems (Zhu et al., 2023; Sun et al., 2023; Gao et al., 2024). To assess whether LLM-based reranking can also enhance the performance of EALink (Zhang et al., 2023b), a deep learning-based method, we conducted experiments on our realistic evaluation dataset, including 20 projects. As shown in Table 7, with LLM reranking assistance, we observe a Precision@1 improvement of 80.45%, increasing from 14.43% to 26.04%. This demonstrates the LLM’s effectiveness in refining initially imprecise results. Future work can explore the optimization of LLM prompts and apply domain-specific fine-tuning to further improve the refinement step.
Table 7. Comparison of EALink and EALink+LLM
Metric EALink EALink+LLM Improvement
P@1 (Hit@1) 14.43 26.04 $\uparrow$ 80.45
P@10 3.64 3.64 -
Hit@10 30.76 30.76 -
Recall@10 28.06 28.06 -
MRR 20.21 28.55 $\uparrow$ 41.26
NDCG@1 9.10 16.96 $\uparrow$ 86.37
NDCG@10 15.52 18.79 $\uparrow$ 21.07
Note: Results (%) are averaged over 20 projects on our dataset constructed under the Realistic Distribution Setting (RDS).
7.2. Threats to Validity
Threats to Internal Validity. Threats to internal validity refer to errors in our experiments or implementation issues. To avoid implementation errors, we replicated the baseline tool, EALink (Zhang et al., 2023b), using its publicly available code. We ensured that we replicated its previously reported results before extending the experiments. Therefore, the threats to internal validity are minimal.
Threats to Construct Validity. A potential threat to construct validity is the selection of evaluation metrics. We use widely adopted metrics—Precision@$k$, Hit@$k$, MRR, and NDCG@$k$—from prior studies (Zhang et al., 2023b; Dong et al., 2022; Ruan et al., 2019; Mazrae et al., 2021) and information retrieval tasks (Radlinski and Craswell, 2010; Sakai and Kando, 2008). We also included Recall@k, a standard information retrieval evaluation metric, overlooked in prior issue-commit linking work. Consequently, we believe that any threat to construct validity is minimal.
Threats to External Validity. Threats to external validity refer to factors that might limit the generalizability of our findings. Our benchmark is the largest in the literature, which provides confidence that our findings are not specific to only a few projects. One threat is the number of times our experiments were performed. To mitigate the effect of randomness, we repeated our experiments five times. For all metrics, the average results exhibit standard deviations below 1%. Given the stability of the results, repeating the experiments would not yield different findings. As such, there are minimal threats to external validity.
8. Related Work
Traceability Link Recovery: Traceability link recovery methods create links between artifacts such as requirements, design documents, architecture models, and source code. Research has applied classic IR techniques (Borg et al., 2014; Lucia et al., 2007; Gethers et al., 2011; Oliveto et al., 2010; Eaddy et al., 2008). Recent work (Rodriguez et al., 2023; Fuchß et al., 2025) utilized LLMs but also found that their level of effectiveness has still been unable to support practical automatic link recovery (Fuchß et al., 2025; Hayes et al., 2006).
Traditional Approaches for Issue-Commit Linking: Traditional approaches combine heuristics and expert annotation to link commits with bug reports. Bachmann et al. (Bachmann et al., 2010) used interactive heuristic linking. Wu et al. (Wu et al., 2011) filtered candidates by textual similarity, timing, and committer mapping. It also learned optimal thresholds from training data. Nguyen et al. (Nguyen et al., 2012) improved performance by adding code change analysis. Schermann et al. (Schermann et al., 2015) leveraged developer identity, time proximity, and resource overlap. These methods often miss links (Bird et al., 2009).
Machine Learning-Based Approaches: Machine learning methods improve linking accuracy by reducing the need for handcrafted heuristics. Le et al. (Le et al., 2015) enriched commit messages via code summarization. Sun et al. (Sun et al., 2017b) refined feature extraction with non-source documents and code file filtering. Sun et al. (Sun et al., 2017a) used positive-unlabeled learning to address limited labeled data. Rath et al. (Rath et al., 2018) combined process, stakeholder, structural, and textual similarity metrics. Mazrae et al. (Mazrae et al., 2021) incorporated non-textual cues such as authorship, timestamps, and status. Dong et al. (Dong et al., 2022), a semi-supervised framework, tackled data imbalance and sparsity. These approaches were later improved by deep learning.
Deep Learning-Based Approaches: Recent works focused on the use of deep learning. Ruan et al. (Ruan et al., 2019) learned semantic representations with word embeddings and RNNs, while Xie et al. (Xie et al., 2019) combined RNNs with SVMs and a code knowledge graph from ASTs for semantic and code context. Lin et al. (Lin et al., 2021) leveraged a BERT-based framework pre-trained on CodeSearchNet and fine-tuned on small datasets to address data sparsity. Zhu et al. (Zhu et al., 2024) employed deep semi-supervised learning and iteratively retrained its model using pseudo-labels on unlabeled data. Zhang et al. (Zhang et al., 2023b) employed knowledge distillation to improve both accuracy and efficiency. Despite these improvements, we found that the approaches were assessed using evaluations whose realism could be improved.
Replication Studies: Our study found that the data used in evaluations for issue-commit link recovery may have the drawback of an unrealistic distribution of false links. Other studies (Shepperd et al., 2013; Ding et al., 2024; Chakraborty et al., 2024; Liu et al., 2025; Kang et al., 2022) have emphasized the importance of using methods and data that evaluate tools in different settings. While other works (Dong et al., 2022; Lin et al., 2021; Mazrae et al., 2021; Ruan et al., 2019; Zhang et al., 2023b) have investigated issues of data cleanliness and data leakage, our work is the first replication study of issue-commit linkers and highlights the importance of evaluating them using realistic evaluation data that match the real development history.
9. Conclusion and Future Work
In this study, we successfully replicate the strong performance of the state-of-the-art work on issue-commit linking. To investigate it further, we constructed a new benchmark under a proposed Realistic Distribution Setting (RDS) that adaptively constructs false links based on the level of development activity in the same time frame as the ground-truth link. To the best of our knowledge, the benchmark is the largest in the literature, consisting of 9,319 unique issues from 20 open-source projects, with an average of 1,530 false links constructed per issue. We find that the use of an off-the-shelf IR method outperforms the state-of-the-art technique. Building on our findings, we propose EasyLink, a scalable and efficient approach that combines retrieval with an additional step of reranking using an LLM. In terms of average Precision@1, EasyLink outperforms the state-of-the-art approach by more than 4 times.
In the future, we plan to extend EasyLink for other tasks in software traceability, such as mapping features to their implementations and linking requirements to code changes.
10. Data Availability
The implementation and documentation of EasyLink, along with our collected evaluation dataset, have been anonymized and made available at this URL link
https://figshare.com/s/d495f11c4cc5c1c72e68

### v2

@@
 To address the long-standing problem of the semantic gap between issues and commits, EasyLink leverages a large language model to rerank the commits retrieved from the database.
-Under our evaluation, EasyLink achieves an average Precision@1 of 75.91%, improving over the state-of-the-art by over four times.
+Under our evaluation, EasyLink achieves an average Precision@1 of 75.03%, improving over the state-of-the-art by over four times.
 Additionally, this paper provides practical guidelines for advancing research in issue-commit link recovery.
 Issue-Commit Link Recovery, Software Traceability
-Jieke Shi is the corresponding author.
-copyright: acmlicensedjournalyear: 2026conference: IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/2018/06
+Jieke Shi and Ting Zhang are the corresponding authors.
+journalyear: 2026copyright: rightsretainedconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3764530isbn: 979-8-4007-2025-3/26/04
 1. Introduction
@@
 ❷ Another limitation is the unrealistic false link distributions in prior evaluation datasets.
-Some (Lin et al., 2021; Ruan et al., 2019) use evaluation datasets where the number of false and true links is equal, an unrealistic assumption since, in reality, false links far outnumber true links (Dong et al., 2022).
+Some (Lin et al., 2021; Ruan et al., 2019) use evaluation datasets where the number of false and true links is equal, an unrealistic assumption since false links far outnumber true links in reality (Dong et al., 2022).
 Recently, Zhang et al.
@@
 We use LLMs for their strong ability to capture the semantic relationship between issues and commits, bridging the semantic gap, as the most similar commit does not guarantee that it is the fix for the issue (Zhang et al., 2023b; Lin et al., 2021; Dong et al., 2022).
-EasyLink achieves an average Precision@1 of 75.91%, outperforming EALink (Zhang et al., 2023b) by 4$\times$ in our realistic datasets and by 30% in Precision@1 on EALink’s original evaluation setup.
+EasyLink achieves an average Precision@1 of 75.03%, outperforming EALink (Zhang et al., 2023b) by four times in our realistic datasets and by more than 30% in Precision@1 on EALink’s original evaluation setup.
 We also conduct an ablation analysis on each of EasyLink’s two stages.
 Our results demonstrate that advances in IR, such as context-aware dense embeddings (Wang et al., 2020; Song et al., 2020; Neelakantan et al., 2022) and efficient vector search (Douze et al., 2024; Wang et al., 2021), have been largely overlooked in the software traceability literature.
-Notably, even the out-of-the-box use of a vector database achieves a high average Precision@1 of 64.03%.
-Additionally, the strong performance of the reranking step, which improves Precision@1 by 11%, highlights the capability of LLMs to bridge the semantic gap between issues and commits (Rath et al., 2018; Guo et al., 2017).
+Notably, even the out-of-the-box use of a vector database achieves a high average Precision@1 of 61.57%.
+Additionally, the strong performance of the reranking step, which improves Precision@1 by 13.46%, highlights the capability of LLMs to bridge the semantic gap between issues and commits (Rath et al., 2018; Guo et al., 2017).
 Finally, we discuss the lessons learned from our work, such as the need to consider modern IR baselines, for future research on software traceability.
-Our implementation, along with the evaluation dataset, has been anonymized and made available at (Package, 2025).
+Our implementation has been made available at (Package, 2025).
 This paper makes the following contributions:
@@
     A new state-of-the-art for issue-commit linking: We propose EasyLink, which leverages modern IR techniques, including an off-the-shelf database FAISS (Douze et al., 2024), and addresses the problem of semantic gap (Rath et al., 2018; Guo et al., 2017) by prompting an LLM to rerank the retrieved results.
-In our evaluation, EasyLink improves over EALink in average Precision@1 from 14% to 76%.
+In our evaluation, EasyLink improves over EALink in average Precision@1 from 14.43% to 75.03%.
 The rest of the paper is organized as follows.
@@
 Section 6 presents the experimental results.
-Section 7 discusses the lessons learned and threats to validity.
+Section 7 revisits the classical baseline, discusses the lessons learned and threats to validity.
 Section 8 reviews related work.
-Section 9 concludes the paper.
-Finally,  Section 10 presents the details of data availability.
+Finally,  Section 9 concludes the paper.
 2. Background
 2.1. Issue-Commit Linking Recovery
-An example of an issue, the incorrectly top-ranked commit, and the correct commit.
-The input issue, with summary and description, is shown below.
-Issue ID: NETBEANS-803 Summary: nb-javac 11 upgrade in NetBeans Description: Should cover below tasks in NetBeans - nb-javac 11 testing : Run tests for modules java.completion, java.editor, java.editor.base, java.hints, java.source, java.source.base, lib.nbjavac Update libs.javacimpl and libs.javacapi jars, upload updated nb-javac jars Upload nb-javac module jars in update center   The top-ranked commit message in the initial results is shown below but is incorrect.
+An example of an issue, the incorrectly top-ranked commit, and the correct commit. The input issue, with summary and description, is shown below.
+Issue ID: NETBEANS-803
+Summary:
+nb-javac 11 upgrade in NetBeans
+Description:
+Should cover below tasks in NetBeans - nb-javac 11 testing :
+• Run tests for modules java.completion, java.editor,
+ java.editor.base, java.hints, java.source,
+ java.source.base, lib.nbjavac
+• Update libs.javacimpl and libs.javacapi jars, upload updated
+ nb-javac jars
+• Upload nb-javac module jars in update center
+  The top-ranked commit message in the initial results is shown below but is incorrect.
 Commit ID: 4fd115aeae3b8423de9ced22d52914e60a1c5800.
-Updation for external nb-javac jar in libs.javacapi and libs.javaimpl modules with nb-javac jar for jdk-12   The correct commit message is shown below.
-Commit ID: 3055661e4dd1c7d587012c917cbec31b27ae9e34 Uptake nb-javac 11 jars for java tests runtime
+Updation for external nb-javac jar in libs.javacapi and
+libs.javaimpl modules with nb-javac jar for jdk-12
+  The correct commit message is shown below.
+Commit ID: 3055661e4dd1c7d587012c917cbec31b27ae9e34
+Uptake nb-javac 11 jars for java tests runtime
 Figure 1.
@@
 We selected datasets from four recent studies, covering 20 open-source software projects over a span of 20 years.
+These datasets use explicit issue tags (e.g., “#123”, “JIRA-456”) in commit messages to construct the true links.
 We include the same issues and true links from these datasets, resulting in a total of 9,319 issues in the benchmark.
@@
 12
-[…]
+13    candidate_commits $\leftarrow$ { commit $\in$ commits_pool $\mid$
 14        commit.time $\geq$ issue.create_time and
@@
 17
-[…]
+18    false_links $\leftarrow$ {(issue, commit, label=0) $\mid$
 19        commit $\in$ candidate_commits }
@@
 By linking $C_{j}$ to $I_{i}$, we generate the false link $f_{i}=\{I_{i},C_{j}\}$.
-[…]
+(1) $\text{is\_candidate}(I_{i},C_{j})$ $=created(I_{i})\leq committed(C_{j})$
 $\land\;committed(C_{j})\leq created(I_{i})+\epsilon,$
-[…]
-[…]
+$\epsilon$ $=\text{one year.}$
+Algorithm 1 details the construction of the evaluation dataset.
+We first split the entire ground-truth dataset into training and test sets following a 4:1 ratio (Zhang et al., 2023b; Dong et al., 2022; Ruan et al., 2019).
+The inputs to the algorithm are the true links from the test set and the commits pool for the project, which is prepared using the process described in Subsection 3.2.
+First, following the method used by EALink (Zhang et al., 2023b), we randomly sample up to 1,000 unique issue IDs from the test set, forming the list $\mathit{sampled\_issue\_id\_list}$.
+If fewer than 1,000 unique issues are available, all issues are included.
+Then, for each $\mathit{issue\_id}$ in this list, we extract the true links for that issue from the test set.
+We use the term “links” (plural) because an issue may be linked to more than one commit (Zhang et al., 2023b).
+In line 6 of the algorithm, we extract only the issue information for the issue being processed (including issue ID, summary, description, etc.), denoted as $\mathit{issue}$.
+Next, as shown in lines 7 to 10 of the algorithm, we filter the $\mathit{candidate\_commits}$ to include all commits whose creation time is no earlier than the issue’s creation time and strictly earlier than one year after the issue’s creation time.
+Additionally, the commit should not be the fix commit of the issue.
+We then construct the $\mathit{false\_links}$ by pairing the $\mathit{issue}$ with each commit in $\mathit{candidate\_commits}$ and labeling the pair as 0.
+Finally, we append both the $\mathit{true\_links}$ and the $\mathit{false\_links}$ for the issue to the evaluation dataset and then proceed to process the next issue.
+Note that this algorithm is executed separately for each project, as each project has its own test dataset and commits pool.
+Following prior work (Zhang et al., 2023b; Ruan et al., 2019), issue tags (e.g., “#123”, “JIRA-456”) in the issue and commit text are removed to prevent potential data leakage.
 Table 1. Statistics of the benchmark
@@
 Isis 24945 652 1265.91
-Netbeans 10502 159 1279.68
+Netbeans 10501 159 1279.68
 Derby 8040 43 363.93
@@
 Pig 3675 45 197.44
-Flask 5353 151 357.88
+Flask $53$ 151 357.88
 Keras 11248 111 454.23
@@
 We believe that switching FAISS with another vector database with comparable capability would produce similar results, as shown in  Subsection 6.3,
-To retrieve commits from the vector database, we compute the squared L2 (Euclidean) distance between the query embedding and the stored embeddings, as we use the vector database through LangChain (Chase, 2022), where L2 distance is used by default.
+To retrieve commits from the vector database, we compute the cosine similarity score, following prior studies (Zhang et al., 2023b; Ruan et al., 2019; Rath et al., 2018).
 The computation is given by the following equation:
-(3) $distance(\mathbf{q},\mathbf{e}_{i})=|\mathbf{q}-\mathbf{e}_{i}|_{2}^{2}$
-where $e_{i}$ is the element of the commit message embeddings $\mathbf{E}=\langle\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{n}\rangle$ that are indexed by the vector database for efficient access, and $\mathbf{q}$ is an issue embedding.
-A shorter distance between the query issue and the commit message indicates greater similarity, enabling the ranking of commit messages and the generation of a list of candidate commits.
+(3) $\operatorname{sim}(\mathbf{q},\mathbf{e}_{i})\;=\;\frac{\mathbf{q}\cdot\mathbf{e}_{i}}{\lVert\mathbf{q}\rVert_{2}\,\lVert\mathbf{e}_{i}\rVert_{2}}$
+where $\mathbf{e}_{i}$ is the $i$-th element of the commit message embeddings $\mathbf{E}=\langle\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{n}\rangle$ that are indexed by the vector database for efficient access, and $\mathbf{q}$ is the issue embedding.
+A higher score between the query issue and the commit message indicates greater similarity, enabling the ranking of commit messages and the generation of a list of candidate commits.
 4.2. LLM-Assisted Reranking
@@
 While it is possible to rerank all fetched commits for higher precision, this requires a longer prompt for the LLM, requiring more computation resources.
-For the value of $k$, we select $k$=10 for its balance of precision and efficiency.
+For the value of $k$, we select kk=10 for its balance of precision and efficiency.
 We later show that this allows a high Precision@1 without incurring a high cost.
@@
   • Precision@k evaluates the proportion of relevant commits (i.e., commits that belong to the correct issue-commit links) within the top $k$ retrieved results:
-[…]
+    (4) $\text{Precision@k}=\frac{1}{|Q|}\sum_{i\in Q}\frac{\text{Rel}_{i}}{k},$
     where $Q$ is the query set, $|Q|$ its size, and $\text{Rel}_{i}$ the number of correctly linked commits in the top $k$ results for query $i$.
   • Hit@k measures the likelihood that at least one correct commit appears within the top $k$ retrieved results:
-[…]
+    (5) $\text{Hit@k}=\frac{1}{|Q|}\sum_{i}\mathbb{I}(\text{Rank}_{i}\leq k),$
     where $\mathbb{I}(\cdot)$ returns 1 if the highest-ranked relevant commit for query $i$ is within the top $k$, and 0 otherwise. Hit@1 is equivalent to Precision@1.
   • Recall@k evaluates the proportion of relevant commits retrieved within the top $k$ results:
-[…]
+    (6) $\text{Recall@k}=\frac{1}{|Q|}\sum_{i\in Q}\frac{\text{Rel}_{i}}{\text{TotalRel}_{i}},$
     where $Q$ is the query set, $|Q|$ its size, $\text{Rel}_{i}$ the number of retrieved relevant commits, and $\text{TotalRel}_{i}$ the total relevant commits for query $i$.
@@
   • NDCG@k (Normalized Discounted Cumulative Gain) assesses how well relevant commits are ranked within the top $k$ retrieved results:
-[…]
+    (8) $\text{NDCG@k}=\frac{1}{Z_{k}}\sum_{i=1}^{k}\frac{2^{r_{i}}-1}{\log_{2}(i+1)},$
     where $Z_{k}$ is a normalization factor ensuring the ideal ranking achieves a value of 1.
@@
 (Zhang et al., 2023b), which had a fixed number of 99 false links per issue, our evaluation generates more false links for repositories with a larger number of commits.
-This results in an average number of false links per issue of 1529.78.
+This results in an average number of false links per issue of 1,530.
 In the second column of Table 2, we present the results of running the tool on our evaluation dataset, which shows a significant performance drop compared to the original evaluation dataset.
@@
 EALink Vector DB EasyLink
-P@1 (Hit@1) 53.67 81.59 89.58($\uparrow 66.90\%$)
-P@10 8.94 11.08 11.08($\uparrow 23.93\%$)
-Hit@10 72.56 92.44 92.44($\uparrow 27.39\%$)
-Recall@10 69.03 89.43 89.43($\uparrow 29.55\%$)
-MRR 60.13 85.57 90.80($\uparrow 51.00\%$)
-NDCG@1 33.86 51.48 56.52($\uparrow 66.92\%$)
-NDCG@10 41.33 56.07 58.03($\uparrow 40.41\%$)
+P@1 (Hit@1) 53.67 81.83 90.04($\uparrow 67.77\%$)
+P@10 8.94 11.14 11.14($\uparrow 24.61\%$)
+Hit@10 72.56 92.89 92.89($\uparrow 28.02\%$)
+Recall@10 69.03 89.93 89.93($\uparrow 30.28\%$)
+MRR 60.13 85.87 91.38($\uparrow 51.97\%$)
+NDCG@1 33.86 51.63 56.81($\uparrow 67.78\%$)
+NDCG@10 41.33 56.31 58.07($\uparrow 40.50\%$)
 Note: The parentheses in the EasyLink column show the improvements over EALink. Bold numbers indicate the highest performance. All values are in %.
@@
 Metric EALink VSM Vector DB EasyLink
-P@1 (Hit@1) 14.43 46.59 64.03 75.91 ($\uparrow 426.1\%$)
-P@10 3.64 8.70 10.82 10.82 ($\uparrow 197.3\%$)
-Hit@10 30.76 70.94 85.35 85.35 ($\uparrow 177.5\%$)
-Recall@10 28.06 67.22 81.48 81.48 ($\uparrow 190.4\%$)
-MRR 20.21 55.10 71.59 79.92 ($\uparrow 295.4\%$)
-NDCG@1 9.10 29.39 40.40 47.89 ($\uparrow 426.3\%$)
-NDCG@10 15.52 39.45 49.37 52.34 ($\uparrow 237.2\%$)
+P@1 (Hit@1) 14.43 46.59 61.57 75.03 ($\uparrow 420.0\%$)
+P@10 3.64 8.70 10.09 10.09 ($\uparrow 177.2\%$)
+Hit@10 30.76 70.94 83.38 83.38 ($\uparrow 171.1\%$)
+Recall@10 28.06 67.22 78.84 78.84 ($\uparrow 181.0\%$)
+MRR 20.21 55.10 69.21 78.52 ($\uparrow 288.5\%$)
+NDCG@1 9.10 29.39 38.85 46.86 ($\uparrow 415.0\%$)
+NDCG@10 15.52 39.45 47.98 51.08 ($\uparrow 229.1\%$)
 Note: Results are averaged over 20 projects (%).
@@
 To ensure a fair comparison, we first evaluated the vector database and EasyLink on the dataset constructed using the EALink method (Zhang et al., 2023b), which generates 99 false links per issue and includes the six projects provided by EALink.
-The results, shown in Table 3, indicate that EasyLink achieves a Precision@1 of 89.58%, significantly outperforming EALink’s 53.67%.
-The vector database method also outperforms EALink, achieving a Precision@1 of 81.59%.
+The results, shown in Table 3, indicate that EasyLink achieves a Precision@1 of 90.04%, significantly outperforming EALink’s 53.67%.
+The vector database method also outperforms EALink, achieving a Precision@1 of 81.83%.
 Furthermore, we compared them using our more realistic evaluation dataset.
-As presented in Table 4, the vector database approach achieves an average Precision@1 of 64.03%, a significant improvement over EALink’s 14.43%.
-With LLM-assisted reranking, EasyLink further enhances Precision@1 by an average of 11.88%, reaching 75.91%, representing a 426.1% increase compared to EALink.
-Following the Mann-Whitney U test (Mann and Whitney, 1947), the improvement of EasyLink over EALink in every metric is statistically significant (p-value $<$ 0.01) and exhibits a large effect size (Cohen, 2013) (Cohen’s D $>$ 0.8).
+As presented in Table 4, the vector database approach achieves an average Precision@1 of 61.57%, a significant improvement over EALink’s 14.43%.
+With LLM-assisted reranking, EasyLink further enhances Precision@1 by an average of 13.46%, reaching 75.03%, representing a 420.0% increase compared to EALink.
+Following the Mann-Whitney U test (Mann and Whitney, 1947), the improvement of EasyLink over EALink in every metric is statistically significant (p-value << 0.01) and exhibits a large effect size (Cohen, 2013) (Cohen’s D >> 0.8).
 These results demonstrate that EasyLink is effective in issue-commit link recovery and that LLMs can be effectively leveraged to enhance performance.
@@
 Train (Total / Per Link) 100.68h / 4.39s N/A N/A
-Test (Total / Per Issue) 17.78h / 6.87s 2.60h / 1.01s 17.76h / 6.86s
+Test (Total / Per Issue) 17.78h / 6.87s 2.26h / 0.87s 14.02h / 5.42s
 Note: “Total” indicates the overall time for all 20 projects.
@@
 However, approaches using a vector database do not require training on a specific issue-commit link dataset, making them easier to use and more practical with less human effort.
-Regarding testing time, EALink requires a total of 17.78 hours for the 20 projects, averaging 6.87 seconds per issue, while the vector database approach takes only 2.60 hours in total, averaging 1.01 seconds per issue.
-These differences are statistically significant according to the Mann-Whitney U test (Mann and Whitney, 1947) (p-value $<$ 0.01) and demonstrate a large effect size (Cohen’s D $>$ 0.8).
-Additionally, the LLM-assisted method (i.e., EasyLink) takes 17.76 hours for testing, which is slightly faster than EALink but achieves a significantly higher performance.
+Regarding testing time, EALink requires a total of 17.78 hours for the 20 projects, averaging 6.87 seconds per issue, while the vector database approach takes only 2.26 hours in total, averaging 0.87 seconds per issue.
+These differences are statistically significant according to the Mann-Whitney U test (Mann and Whitney, 1947) (p-value << 0.01) and demonstrate a large effect size (Cohen’s D >> 0.8).
+Additionally, the LLM-assisted method (i.e., EasyLink) takes 14.02 hours for testing, which is faster than EALink and achieves a significantly higher performance.
 These results demonstrate that the use of vector database is an efficient solution.
@@
 By incorporating an LLM reranking step, EasyLink achieves a high precision while eliminating the need for training models.
-Answer to RQ2: Our results show that on the realistic evaluation dataset, EasyLink achieves a Precision@1 of 75.91%, outperforming EALink (which achieves 14.43%) by 426.1%.
+In practice, each issue requires about 758 input tokens and 266 output tokens, leading to an estimated cost of only \$0.009 per issue when using GPT-4o.
+Answer to RQ2: Our results show that on the realistic evaluation dataset, EasyLink achieves a Precision@1 of 75.03%, outperforming EALink (which achieves 14.43%) by 420.0%.
 Moreover, using the vector database-based method eliminates the need for pretraining on task-specific data, making it a more efficient solution.
@@
 MiniLM MPNet OpenAI
-P@1 (Hit@1) 64.03 61.22 72.16
-P@10 10.82 10.37 11.09
-Hit@10 85.35 82.39 88.64
-Recall@10 81.48 78.58 84.73
-MRR 71.59 69.02 78.40
-NDCG@1 40.40 38.65 45.53
-NDCG@10 49.37 47.62 52.67
-Test Time (hour) 2.60h 3.10h 29.61h
+P@1 (Hit@1) 61.57 62.73 70.10
+P@10 10.09 10.18 10.62
+Hit@10 83.38 83.00 87.08
+Recall@10 78.84 78.40 82.91
+MRR 69.21 70.39 76.51
+NDCG@1 38.85 39.58 44.22
+NDCG@10 47.98 48.32 51.56
+Test Time (hour) 2.26h 3.22h 28.72h
 Model Size 22.7M params 109M params Cloud-based
@@
 Table 6 presents the results comparing three embedding models.
-The comparison between MiniLM (used in EasyLink) and MPNet shows that the model we selected in our tool requires less time while achieving relatively better performance.
-Although OpenAI embeddings provide better performance—with Precision@1 increasing from 64.03% to 72.16% (an improvement of 8.13%)—they require 29.61 hours to complete the experiment, approximately 11 times longer than the MiniLM model.
+The comparison between MiniLM (used in EasyLink) and MPNet shows that the model we selected in our tool requires less time while achieving comparable performance.
+Although OpenAI embeddings provide better performance—with Precision@1 increasing from 61.57% to 70.10% (an improvement of 8.53%)—they require 28.72 hours to complete the experiment, approximately 11 times longer than the MiniLM model.
 Given the high computational cost, MiniLM is a more practical option.
@@
 We conducted this experiment across 20 projects using the realistic evaluation dataset.
-The results in Figure 4 show that increasing $k$—meaning reranking a larger set of top results—generally leads to better performance, specifically higher Precision@1 and NDCG@1.
+The results in Figure 4 show that increasing kk—meaning reranking a larger set of top results—generally leads to better performance, specifically higher Precision@1 and NDCG@1.
 However, this improvement comes with a significant increase in time cost and requires processing more input tokens, leading to a higher cost in invoking OpenAI API.
@@
 7. Discussion
+7.1. Revisiting Classical Baselines
+In addition to recent learning-based methods, it is important to include classical baselines when revisiting research progress on issue-commit link recovery.
+We therefore evaluate ReLink (Wu et al., 2011), an approach that extends beyond the basic VSM, using our constructed dataset.
+Since ReLink outputs positive issue-commit links rather than ranked candidates, we aligned the evaluation by treating EasyLink’s top-ranked commit per issue as its predicted link and computed precision, recall, and F1 score accordingly.
+The average results across 20 projects show that ReLink achieves 10.92% precision, 12.59% recall, and 10.79% F1 score, whereas EasyLink reaches 75.03% precision, 50.75% recall, and 58.75% F1 score, significantly outperforming ReLink.
+Unlike ReLink, which relies on explicit issue tags in commit messages (removed in our setup) and operates at the set level, EasyLink leverages semantic relevance and ranks candidates per issue, making it more robust in large and noisy candidate pools.
 7.2. Lessons learned and Implications
@@
 A potential threat to construct validity is the selection of evaluation metrics.
-We use widely adopted metrics—Precision@$k$, Hit@$k$, MRR, and NDCG@$k$—from prior studies (Zhang et al., 2023b; Dong et al., 2022; Ruan et al., 2019; Mazrae et al., 2021) and information retrieval tasks (Radlinski and Craswell, 2010; Sakai and Kando, 2008).
+We use widely adopted metrics—Precision@$k$, Hit@$k$, MRR, and NDCG@kk—from prior studies (Zhang et al., 2023b; Dong et al., 2022; Ruan et al., 2019; Mazrae et al., 2021) and information retrieval tasks (Radlinski and Craswell, 2010; Sakai and Kando, 2008).
 We also included Recall@k, a standard information retrieval evaluation metric, overlooked in prior issue-commit linking work.
@@
 Traceability Link Recovery: Traceability link recovery methods create links between artifacts such as requirements, design documents, architecture models, and source code.
-Research has applied classic IR techniques (Borg et al., 2014; Lucia et al., 2007; Gethers et al., 2011; Oliveto et al., 2010; Eaddy et al., 2008).
+Research has applied classic IR techniques (Borg et al., 2014; Lucia et al., 2007; Gethers et al., 2011; Oliveto et al., 2010; Eaddy et al., 2008; Yang et al., 2021).
 Recent work (Rodriguez et al., 2023; Fuchß et al., 2025) utilized LLMs but also found that their level of effectiveness has still been unable to support practical automatic link recovery (Fuchß et al., 2025; Hayes et al., 2006).
@@
 Replication Studies: Our study found that the data used in evaluations for issue-commit link recovery may have the drawback of an unrealistic distribution of false links.
-Other studies (Shepperd et al., 2013; Ding et al., 2024; Chakraborty et al., 2024; Liu et al., 2025; Kang et al., 2022) have emphasized the importance of using methods and data that evaluate tools in different settings.
-While other works (Dong et al., 2022; Lin et al., 2021; Mazrae et al., 2021; Ruan et al., 2019; Zhang et al., 2023b) have investigated issues of data cleanliness and data leakage, our work is the first replication study of issue-commit linkers and highlights the importance of evaluating them using realistic evaluation data that match the real development history.
+Some other studies (Shepperd et al., 2013; Ding et al., 2025; Chakraborty et al., 2024; Liu et al., 2024; Kang et al., 2022; Shi et al., 2025) have also emphasized the importance of using methods and data that evaluate tools in different settings.
+While some works (Dong et al., 2022; Lin et al., 2021; Mazrae et al., 2021; Ruan et al., 2019; Zhang et al., 2023b) have investigated issues of data cleanliness and data leakage, our work is the first replication study of issue-commit linkers and highlights the importance of evaluating them using realistic evaluation data that match the real development history.
 9. Conclusion and Future Work
@@
 Building on our findings, we propose EasyLink, a scalable and efficient approach that combines retrieval with an additional step of reranking using an LLM.
-In terms of average Precision@1, EasyLink outperforms the state-of-the-art approach by more than 4 times.
+In terms of average Precision@1, EasyLink outperforms the state-of-the-art approach by more than four times.
 In the future, we plan to extend EasyLink for other tasks in software traceability, such as mapping features to their implementations and linking requirements to code changes.
-10. Data Availability
-The implementation and documentation of EasyLink, along with our collected evaluation dataset, have been anonymized and made available at this URL link
-https://figshare.com/s/d495f11c4cc5c1c72e68
+Acknowledgements.
+This research / project is supported by the National Research Foundation, Singapore, and the Smart Nation Group under the Smart Nation Group’s Translational R&D Grant (Award No.
+TRANS2023-TGC02).
+Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore or the Smart Nation Group.